Compound interest. How many years will it take for an investment of 600€ to grow to 1500€ if the annual interest rate is 9% and the interest is compounded annually?
Answer #1

Future value = Present value * ( 1 + rate ) n

1500 = 600 *( 1+9%) n

n = 10.63 years or 11 years approx

Alternatively using factor tables,

Future value of annuity due = 1500 / 600 = 2.5

Looking at 9% interest rate, no of years = 11 years (approx.).